OpenStudy (anonymous):

Which of the following equations is an example of direct variation? y = x - 2 y = 2(x +5) x - 3y = 0 x = y + 3

OpenStudy (anonymous):

It 's supposed to look like y=kx, if it is y=kx+b it is indirect.

OpenStudy (anonymous):

So it could be the second choice?

OpenStudy (cggurumanjunath):

x - 3y = 0 as per @GABI_J

OpenStudy (anonymous):

Yes, Of course.

OpenStudy (cggurumanjunath):

x - 3y = 0 IS THE ANSWER @hottiewithabody

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if you rearrange it, hottiewithabody , you will see that. add 3y to both sides, and divide both sides by 3, x's coefficient will be the direct variation.